OpenStudy (anonymous):

(1- cos^2x) (1+ cot^2x)=1

OpenStudy (anonymous):

\[(1- \cos^2x) (1+ \cot^2x)=\sin^2x(1+\frac{ \cos^2x}{\sin^2x})=\sin^2x+\cos^2x=1\]

OpenStudy (anonymous):

Thank you so much ....how do you convert cos too sin in these problems ?

OpenStudy (anonymous):

Recall that \[\sin^2x+\cos^2x=1 \]

OpenStudy (anonymous):

so\[1-\sin^2x=\cos^2x~~and~~1-\cos^2x=\sin^2x\]

OpenStudy (anonymous):

okay i see now. may you help me with a few more problems

OpenStudy (anonymous):

Sorry, time for me to go to bed. Gotta get up in the morning. Do math every day.

OpenStudy (anonymous):

aww man ok